OpenStudy (anonymous):

Solve the following system using the substitution method. x + 3y = –5 5x + 6y = –19

OpenStudy (anonymous):

x=-3y-5 substitute in 2nd eqn 5(-3y-5)+6y=-19 -15y-25+6y=-19 -9y=6 y=-2/3 put back y to get x x=(-3*-2/3)-5=-3 solution x=-3, y=-2/3